OpenStudy (anonymous):

Someone Please Help! Solve a system of linear equations using substitution, elimination OR substitution! 3x - y = -137 y = 2x + 99

OpenStudy (anonymous):

what is x?

OpenStudy (anonymous):

3x - y = -137 ...........(1) y = 2x + 99 ..............(2) substitute (2) into (1) 3x -(2x+99) = -137 try solving it now

OpenStudy (anonymous):

good reply!

OpenStudy (nubeer):

put your second equation in first one 3x- ( 2x+99) = -137 find x now.

OpenStudy (anonymous):

i got y = 23 for substituting. Now i have 2 more equations that i have to find for graphing and eliminating, can someone help me witht hose 2??

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

what r they?

OpenStudy (anonymous):

THANK YOUU!! 2.) 2x + y = -1 x + 4y = -10 3.) 2x + y = -1 3x - y = 16 And i have to solve it by substituting and the other by graphing:\

OpenStudy (anonymous):

sorry, i meant *eliminating* and graphing!

OpenStudy (anonymous):

ok mmmmmmmmhhhhhhhhhhhhhhhhhhh... i am not sure how much help that i will be since i am 10 years old :/

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Oh myyyy haha, at least you know more than me. I'm 17! Your smart. But if you can just try? or start me off and i can try to figure out the rest by myself

OpenStudy (anonymous):

ok i will try... gtg sorry :((

OpenStudy (anonymous):

kay, bye. thanks tho

OpenStudy (anonymous):

for 2) by elimination 2x+y=-1...........(1) x+4y = -10.......(2) (2) x 2 2x+8y = -20......(3) (3) -(1) 7y = -19 k solve it now

OpenStudy (anonymous):

Wait, 2x + 8y = -20, i can't solve it no more, can i ? idk, i'm confused! how do you get (3) - (1) ?

OpenStudy (anonymous):

(3) - (1) as in equation (3) - equation(1) to eliminate the 2x which is common in both equations

OpenStudy (anonymous):

basically i manipulate (2) so that the x component looks like (1). I then label the manipulated (1) as a new equation (3). (3) and (1) have 2x. so if i take (3) - (1) i will eliminate the 2x and get left with 7y = -19

OpenStudy (anonymous):

2x + y = -1......(1) 2x + 8y = -20...(3) (3) - (1) 2x - 2x +8y - y = -20 - (-1) 7y = -19

OpenStudy (anonymous):

Ohh! okay okay, i understand.. so y = -2.714 ?

OpenStudy (anonymous):

@irkiz thank you anyways